OpenStudy (anonymous):

Choose the system of inequalities that matches the graph below. A. 3x + 5y > –10 x – 3y > –12 B. 3x + 5y > –10 x – 3y < –12 C. 3x + 5y greater than or equal to –10 x – 3y greater than or equal to –12 D. 3x + 5y greater than or equal to –10 x – 3y lesser than or equal to –12

OpenStudy (luigi0210):

There's no shading?

OpenStudy (anonymous):

oh sorry wrong graph hold on

OpenStudy (anonymous):

any ideas @mebs ?

OpenStudy (luigi0210):

Well your lines are solid, so that means that the graph must have a "or equal to"

OpenStudy (anonymous):

so either C or D

OpenStudy (luigi0210):

Yup. Any guesses?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

?

OpenStudy (anonymous):

From the other graph given haha yea looking at... I would say that it probably can't already say it is not A or B because this is linear graph. From C and D. C \[y \ge -3/5x -2\] \[y \le 1/3x +4\] for D \[y \ge -3/5x -2\] \[y \ge 1/3x +4\] From the two choices the better would be D because you can just check by solving for x or y.

OpenStudy (luigi0210):

Hm, sure what Mebs said..

OpenStudy (anonymous):

where are you getting all these questions from @ pamelasaucedo .....are you still going to school no summer breaks